OpenStudy (anonymous):

Please help solve, I'm very confused. (180)(0.05)P+600=(375)(0.04)P

OpenStudy (texaschic101):

(180)(0.05)P + 600 = (375)(0.04)P 9p + 600 = 15p 600 = 15p - 9p 600 = 6p 600/6 = p 100 = p

OpenStudy (anonymous):

Thank you!

OpenStudy (texaschic101):

no problem :)